OpenStudy (anonymous):

A cone has a volume of 5 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of? a. 15 in3 b. 20 in3 c. 25 in3 d. 30 in3

OpenStudy (anonymous):

@Directrix ?cx

Directrix (directrix):

Volume of Cone: 1/3 Bh where B is the area of the base 1/3 Bh = 5 Bh = 15 B = 15/h The cylinder has the same base area as the cone. Volume of Cylinder: V = B*h B = V/h and B = 15/h The height of the cone and the cylinder are the same. V/h = 15/h Volume of Cylinder = 15. @lowcard2

OpenStudy (anonymous):

Thank you, i understand cx

OpenStudy (anonymous):

what do you mean?